# An ordering for the strength of functional dependence

### v1

@@ 4 Dependence measures generated by $\preccurlyeq_{ccx}$
    and the version of $\preccurlyeq_{ccx}$ for cumulative probabilities (see Remark 1.3(b)), it follows that also $\nu$ is increasing in the conditional convex order.
@@
For a second class of dependence measures generated by the conditional convex order, we use the characterization of $\preccurlyeq_{ccx}$ via the bivariate concordance order in Theorem 1.6(ii).
@@
Therefore, we denote by $\mathcal{R}^{\uparrow}:=\{(T,S)\mid T,S\sim U(0,1),T\uparrow_{st}S\}$ the class of bivariate SI random vectors with marginals that are uniform on $(0,1)$. Let $\mu$ be a functional on $\mathcal{R}^{\uparrow}$ that is normalized and increasing in the concordance order in the sense that $\mu(T,S)=0$ if and only if $T$ and $S$ are independent, $\mu(T,S)=1$ if and only if $T$ and $S$ are comonotone, and $(T,S)\leq_{c}(T^{\prime},S^{\prime})$ implies $\mu(T,S)\leq\mu(T^{\prime},S^{\prime})$. We say that $\mu$ is strictly increasing on $\mathcal{R}^{\uparrow}$ if $(T,S)\leq_{c}(T^{\prime},S^{\prime})$ with $(T,S)\not\stackrel{{\scriptstyle\mathrm{d}}}{{=}}(T^{\prime},S^{\prime})$ implies $\mu(T,S)<\mu(T^{\prime},S^{\prime})$. For a multivariate random vector $(Y,\mathbb{X})\in\mathcal{R}_{[0,1]}$ and for $V,U$ i.i.d. uniform on $(0,1)$, we define
$\mu(Y,\mathbb{X}):=\mu(F_{Y}\circ q_{Y|\mathbb{X}}^{\uparrow U}(V),U),$ (26)
@@
where $(F_{Y}\circ q_{Y|\mathbb{X}}^{\uparrow U}(V),U)$ is a reduced random vector associated with $(Y,\mathbb{X})$. Since $(Y,\mathbb{X})\in\mathcal{R}_{[0,1]}$, we have that $F_{Y}$ is continuous. Hence, the reduced random vector $(F_{Y}\circ q_{Y|\mathbb{X}}^{\uparrow U}(V),U)$ has uniform marginals. Since it is also SI, it is in $\mathcal{R}^{\uparrow}$, and $\mu(Y,\mathbb{X})$ in (26) is well-defined. Recall the identity $(Y,\mathbb{X})=_{ccx}(F_{Y}\circ q_{Y|\mathbb{X}}^{\uparrow U}(V),U)$ due to Proposition 2.7(iii) and the distributional invariance of the conditional convex order. Consequently, the functional $\mu$ in (26) depends only on the strength of functional dependence of $Y$ on $\mathbb{X}$ in the sense of the conditional convex order. Equivalently, by Theorem 1.6, $\mu$ in (26) depends only on the strength of positive dependence of $(F_{Y}\circ q_{Y|\mathbb{X}}^{\uparrow U}(V),U)$ in the sense of the concordance order.
@@
The functional $\mu$ in (26) defines a dependence measure with the following properties:
@@
    $(Y,\mathbb{X})\preccurlyeq_{ccx}(Y^{\prime},\mathbb{X}^{\prime})$ implies $\mu(Y,\mathbb{X})\leq\mu(Y^{\prime},\mathbb{X}^{\prime})$,
@@
    $\mu(Y,\mathbb{X})=0$ if and only if $\mathbb{X}$ and $Y$ are independent,
@@
    $\mu(Y,\mathbb{X})=1$ if and only if $Y$ perfectly depends on $\mathbb{X}$,
@@
    $\mu(Y,\mathbb{X})\leq\mu(Y,(\mathbb{X},\mathbb{Z}))$,
  5. (v)
    If, additionally, $\mu$ is strictly increasing on $\mathcal{R}^{\uparrow}$, then $\mu(Y,\mathbb{X})=\mu(Y,(\mathbb{X},\mathbb{Z}))$ if and only if $Y$ and $\mathbb{Z}$ are conditionally independent given $\mathbb{X}$.
@@
Our approach on rearranged dependence measures in Theorem 4.4 is studied in [62] for bivariate random vectors with continuous marginal distribution functions. Various examples of measures of concordance that are increasing on $\mathcal{C}^{\uparrow}$, such as Kendall’s $\tau$ and Gini’s $\gamma$, or even strictly increasing, such as Spearman’s $\rho$ or the Schweizer-Wolff measure, are given in [62, Section 2.2].
5 Comparing $\preccurlyeq_{ccx}$ with related dependence orderings
In this section, we compare the conditional convex order with known dependence orderings from the literature. First, we study the fundamentally different concepts of the concordance order and the conditional convex order; see Figure 3 and Table 1. As we show in Proposition 5.1, these orders coincide for stochastically increasing distributions. We then discuss related orderings of functional dependence studied in [22, 59] and demonstrate that the conditional convex order provides a more natural and powerful framework for ordering strengths of dependencies.
@@ 5.1 Comparison of $\preccurlyeq_{ccx}$ with the concordance order
where the lower orthant and upper orthant order on the right-hand side of (27) are defined by $P(Y\leq y,X\leq x)\leq P(Y^{\prime}\leq y,X^{\prime}\leq x)$ and $P(Y>y,X>x)\leq P(Y^{\prime}>y,X^{\prime}>x)$ for all $x,y\in\mathbb{R}$, respectively, see e.g. [32, Definition 2.4]. As a direct consequence of the definition, $(Y^{\prime},X^{\prime})\leq_{c}(Y^{\prime},X^{\prime})$ implies $X\stackrel{{\scriptstyle\mathrm{d}}}{{=}}X^{\prime}$ and $Y\stackrel{{\scriptstyle\mathrm{d}}}{{=}}Y^{\prime}$. Note that for equal marginal distributions, the bivariate concordance order is equivalent to the lower (and also to the upper) orthant order; see Proposition B.2.
To relate the concordance order and the conditional convex order, we consider, for simplicity, bivariate random vectors $(V,U)$ and $(V^{\prime},U^{\prime})$ with marginals that are uniform on $(0,1).$ This corresponds to the setting of bivariate copulas. Using the notation $\eta_{V|U}^{v}(t)=P(V\leq v\mid U=t)$ in (6) and disintegration, the concordance order can be characterized by
@@
An extension of the above reasoning to arbitrary marginals yields the following result. It states that the conditional convex order and the concordance order coincide for bivariate SI random vectors from the same Fréchet class.
@@
Figure 3: Visualization of the concordance order (left plot) and the conditional convex order (right plot) in the case of bivariate copulas. Here, $W$ and $M$ denote the lower and upper Fréchet copula and $\Pi$ the independence copula. While $W$ and $M$ are the uniquely determined minimal and maximal elements in concordance order, the conditional convex order has the independence copula as global minimal element and all perfectly dependent copulas as global maximal elements.
@@
Invariance under increasing transformations $(a(Y),X)\leq_{c}(a(Y^{\prime}),X^{\prime}))$ $(Y,\mathbb{X})\preccurlyeq_{ccx}(a(Y^{\prime}),\mathbb{X}^{\prime})$
Invariance under decreasing transformations $(b(Y),X)\leq_{c}(b(Y^{\prime}),X^{\prime}))$ $(b(Y),\mathbb{X})\preccurlyeq_{ccx}(b(Y^{\prime}),\mathbb{X}^{\prime})$
Invariance under bijective transformations $(Y,\mathbb{X})\preccurlyeq_{ccx}(Y^{\prime},h(\mathbb{X}^{\prime}))$
Stability under convergence weak convergence conditional weak convergence
Table 1: Overview of basic properties of the bivariate concordance order and the conditional convex order; AA is a closed subset of $[0,1]$, $a$ and $b$ are strictly increasing/decreasing functions, $h$ is a bijective function

### v2

@@ 4.1 Dependence measures based on conditional distribution functions
    and the version of $\preccurlyeq_{ccx}$ for cumulative probabilities (see Remark 1.3b), it follows that also $\nu$ is increasing in the conditional convex order.
4.2 Optimal transport-based dependence measures
As a second class of dependence measures generated by $\preccurlyeq_{ccx}$, we consider optimal transport-based Wasserstein correlations as studied in [68]. Therefore, let $\nu$ and $\nu^{\prime}$ be distributions on $\mathbb{R}$ and $c\colon\mathbb{R}\times\mathbb{R}\to\mathbb{R}$ be a continuous cost function satisying $c(y,y^{\prime})\geq a(y)+b(y^{\prime})$ for some $a\in L^{1}(\nu)$ and $b\in L^{1}(\nu^{\prime})$. Then Kantorovich’s mass transportation problem consists of finding an optimal coupling between $\nu$ and $\nu^{\prime}$ that attains the minimal costs
$\mathcal{W}_{c}(\nu,\nu^{\prime}):=\inf_{\gamma\in\Pi(\nu,\nu^{\prime})}\int c(y,y^{\prime})\mathsf{\,d}\gamma(y,y^{\prime}),$ (30)
where $\Pi(\nu,\nu^{\prime})$ is the set of couplings between $\nu$ and $\nu^{\prime}$; see [65]. Now, the Wasserstein correlation coefficient in [68] adapted to our setting is defined by
$\overrightarrow{\mathcal{W}}_{c}(Y,\mathbb{X}):=\frac{\int_{\mathbb{R}^{p}}\mathcal{W}_{c}(\pi_{\mathbb{x}},\nu)\mathsf{\,d}\mu(\mathbb{x})}{\int_{\mathbb{R}}\int_{\mathbb{R}}c(y,y^{\prime})\mathsf{\,d}\nu(y)\mathsf{\,d}\nu(y^{\prime})}.$ (31)
Here $\nu$, $\mu$, and $\pi_{\mathbb{x}}$ denote the distribution of $F_{Y}(Y)$, $\mathbb{X}$, and $F_{Y}(Y)\mid\mathbb{X}=\mathbb{x}$, respectively. For submodular [^4] A function $f\colon\mathbb{R}^{2}\to\mathbb{R}$ is said to be submodular if $f(\mathbb{x})+f(\mathbb{y})\geq f(\mathbb{x}\wedge\mathbb{y})+f(\mathbb{x}\vee\mathbb{y})$ for all $\mathbb{x},\mathbb{y}\in\mathbb{R}^{2}$, where $\wedge$ and $\vee$ denote the componentwise minimum and maximum, respectively. cost functions $c$, it is well known that the comonotone coupling $(F_{\pi_{\mathbb{x}}}^{-1}(V),F_{\nu}^{-1}(V))$ attains the minimal cost $W_{c}(\pi_{\mathbb{x}},\nu)$ in the numerator of (31), i.e.
$\mathcal{W}_{c}(\pi_{\mathbb{x}},\nu)=\mathbb{E}c(F_{F_{Y}(Y)|\mathbb{X}=\mathbb{x}}^{-1}(V),F_{F_{Y}(Y)}^{-1}(V));$ (32)
see [49, Theorem 3.1.2]. Important examples of submodular cost functions are convex costs of the type $c(y,y^{\prime})=h(y^{\prime}-y)$ for $h$ convex. These functions include the standard costs $c(y,y^{\prime})=|y^{\prime}-y|^{p}$ for $p\geq 1$. For the following result, we use the characterization ii in Theorem 1.6 to show that, for convex costs, the Wasserstein correlation $\overrightarrow{\mathcal{W}}_{c}(Y,\mathbb{X})$ is $\preccurlyeq_{ccx}$-increasing and satisfies all the desired properties of a dependence measure that quantifies the strength of functional dependence.
Theorem 4.4 (Optimal transport-based dependence measures)
Assume a cost function $c(y,y^{\prime})=h(y^{\prime}-y)$ where $h\colon\mathbb{R}\to\mathbb{R}$ is convex and strictly convex at $0=h(0)$. Then
$(Y,\mathbb{X})\preccurlyeq_{ccx}(Y^{\prime},\mathbb{X}^{\prime})\quad\text{implies}\quad\overrightarrow{\mathcal{W}}_{c}(Y,\mathbb{X})\leq\overrightarrow{\mathcal{W}}_{c}(Y^{\prime},\mathbb{X}^{\prime}).$ (33)
Further, the following statements hold true.
  1. (i)
    $\overrightarrow{\mathcal{W}}_{c}$ attains values only in $[0,1]$.
  2. (ii)
    $\overrightarrow{\mathcal{W}}_{c}$ has the zero-independence and max-functionality property.
  3. (iii)
    $\overrightarrow{\mathcal{W}}_{c}$ satisfies information monotonicity, i.e. $\overrightarrow{\mathcal{W}}_{c}(Y,\mathbb{X})\leq\overrightarrow{\mathcal{W}}_{c}(Y,(\mathbb{X},\mathbb{Z}))$.
  4. (iv)
    If $h$ is strictly convex, then $\overrightarrow{\mathcal{W}}_{c}(Y,\mathbb{X})=\overrightarrow{\mathcal{W}}_{c}(Y,(\mathbb{X},\mathbb{Z}))$ if and only if $Y$ and $\mathbb{Z}$ are conditionally independent given $\mathbb{X}$.
Remark 4.5
  1. (a)
    The function $h$ in Theorem 4.4 can attain negative values. Hence, also the conditional optimal cost $\mathcal{W}_{c}(\pi_{\mathbb{x}},\nu)$ may be negative. However, by convexity of $h$ and $h(0)=0$, the average cost $\int_{\mathbb{R}^{p}}\mathcal{W}_{c}(\mu_{\mathbb{x}},\nu)\mathsf{\,d}\mu(\mathbb{x})$ is non-negative.
  2. (b)
    As for the dependence measures $\xi_{\varphi}$ and $\Lambda_{\varphi}$ in Theorem 4.1, the Wasserstein correlation $\overrightarrow{\mathcal{W}}_{c}$ inherits invariance properties from from the conditional convex order. More precisely, $\overrightarrow{\mathcal{W}}_{c}(Y,\mathbb{X})=\overrightarrow{\mathcal{W}}_{c}(g(Y),{\bf h}(\mathbb{X}))$ for all strictly increasing functions $g$ and bijective functions $\bf{h}$, and $\overrightarrow{\mathcal{W}}_{c}(Y,\mathbb{X})=\overrightarrow{\mathcal{W}}_{c}(F_{Y}(Y),{\bf F}_{\mathbb{X}}(\mathbb{X}))$.
  3. (c)
    The Wasserstein correlation in [68] is defined with respect to a metric cost function. Apart from the metric $d(y,y^{\prime})=|y^{\prime}-y|$, metrics $d$ on $\mathbb{R}$ are generally not submodular (since metrics are subadditive). Consequently, the representation of $\mathcal{W}_{c}$ in terms of conditional comonotonicity in (32) typically does not apply for metric costs $c=d$.
4.3 Rearrangement-based dependence measures
@@
For a third class of dependence measures generated by $\preccurlyeq_{ccx}$, we use the characterization of $\preccurlyeq_{ccx}$ via the bivariate concordance order in Theorem 1.6iii.
@@
Therefore, we denote by $\mathcal{R}^{\uparrow}:=\{(T,S)\mid T,S\sim U(0,1),T\uparrow_{st}S\}$ the class of bivariate SI random vectors with marginals that are uniform on $(0,1)$. Let $\mu$ be a functional on $\mathcal{R}^{\uparrow}$ that is normalized and increasing in the concordance order in the sense that $\mu(T,S)=0$ if and only if $T$ and $S$ are independent, $\mu(T,S)=1$ if and only if $T$ and $S$ are comonotone, and $(T,S)\leq_{c}(T^{\prime},S^{\prime})$ implies $\mu(T,S)\leq\mu(T^{\prime},S^{\prime})$. We say that $\mu$ is strictly increasing on $\mathcal{R}^{\uparrow}$ if $\leq_{c}$ is strict on $\mathcal{R}^{\uparrow}$, i.e. $(T,S)\leq_{c}(T^{\prime},S^{\prime})$ with $(T,S)\not\stackrel{{\scriptstyle\mathrm{d}}}{{=}}(T^{\prime},S^{\prime})$ implies $\mu(T,S)<\mu(T^{\prime},S^{\prime})$. For $(Y,\mathbb{X})\in\mathcal{R}_{[0,1]}$ (recall (6) for the definition of this class) and for $V,U$ i.i.d. uniform on $(0,1)$, we define the rearranged dependence measure $\mathsf{R}_{\mu}$ by
$\mathsf{R}_{\mu}(Y,\mathbb{X}):=\mu(F_{Y}\circ q_{Y;\mathbb{X}}^{\uparrow U}(V),U),$ (34)
@@
where $(F_{Y}\circ q_{Y;\mathbb{X}}^{\uparrow U}(V),U)$ is a reduced random vector associated with $(Y,\mathbb{X})$. Since $(Y,\mathbb{X})\in\mathcal{R}_{[0,1]}$, we have that $F_{Y}$ is continuous. Hence, the reduced random vector $(F_{Y}\circ q_{Y;\mathbb{X}}^{\uparrow U}(V),U)$ has marginals that are uniform on $(0,1)$. Since it is also SI, it is in $\mathcal{R}^{\uparrow}$, and $\mathsf{R}_{\mu}(Y,\mathbb{X})$ in (34) is well-defined. Recall the identity $(Y,\mathbb{X})=_{ccx}(F_{Y}\circ q_{Y;\mathbb{X}}^{\uparrow U}(V),U)$ in (7). Hence, $\mathsf{R}_{\mu}$ depends on the strength of functional dependence of $Y$ on $\mathbb{X}$ in the sense of the conditional convex order. Equivalently, by Theorem 1.6, $\mathsf{R}_{\mu}$ depends on the strength of positive dependence of $(F_{Y}\circ q_{Y;\mathbb{X}}^{\uparrow U}(V),U)$ in the sense of the concordance order.
@@
For the functional $\mathsf{R}_{\mu}$ in (34), we have that
$(Y,\mathbb{X})\preccurlyeq_{ccx}(Y^{\prime},\mathbb{X}^{\prime})\quad\text{implies}\quad\mathsf{R}_{\mu}(Y,\mathbb{X})\leq\mathsf{R}_{\mu}(Y^{\prime},\mathbb{X}^{\prime}).$ (35)
Further, the following statements hold true.
@@
    $\mathsf{R}_{\mu}$ takes values in $[0,1]$.
@@
    $\mathsf{R}_{\mu}$ has the zero-independence and max-functionality property.
@@
    $\mathsf{R}_{\mu}$ satisfies information monotonicty, i.e. $\mathsf{R}_{\mu}(Y,\mathbb{X})\leq\mathsf{R}_{\mu}(Y,(\mathbb{X},\mathbb{Z}))$.
@@
    If, additionally, $\mu$ is strictly increasing on $\mathcal{R}^{\uparrow}$, then $\mathsf{R}_{\mu}(Y,\mathbb{X})=\mathsf{R}_{\mu}(Y,(\mathbb{X},\mathbb{Z}))$ if and only if $Y$ and $\mathbb{Z}$ are conditionally independent given $\mathbb{X}$.
Considering the functional form of $\mathsf{R}_{\mu}$ in (34), we can extend Wasserstein correlations to non-linear functionals by defining
$\mathsf{S}_{\mu}(Y,\mathbb{X})=1-\mu(F_{Y}\circ F_{Y|\mathbb{X}}^{-1}(V),V)$ (36)
for $V\sim U(0,1)$ independent of $\mathbb{X}$. Then the following result is a direct consequence of Theorem 1.6 and Theorem 4.6.
Corollary 4.7 (Dependence measures generated by conditional comonotonicity)
$\mathsf{S}_{\mu}$ in (36) is a dependence measure that has the same properties as $\mathsf{R}_{\mu}$ in Theorem 4.6.
@@
  1. (a)
    Our approach on rearranged dependence measures in Theorem 4.6 is studied in [62] for the case of bivariate random vectors with continuous marginal distribution functions. Various examples of measures of concordance that are increasing on $\mathcal{R}^{\uparrow}$, such as Kendall’s $\tau$ and Gini’s $\gamma$, or even strictly increasing, such as Spearman’s $\rho$ or the Schweizer-Wolff measure, are given in [62, Section 2.2].
  2. (b)
    For the functional $\mathsf{S}_{\mu}$ in (36), note that the bivariate random vector $(F_{Y}\circ F_{Y|\mathbb{X}}^{-1}(V),V)$ is $\leq_{c}$-decreasing in $\preccurlyeq_{ccx}$ by equivalence of Theorem 1.6i and $i$. If $\mu$ is a linear functional of the form $\mathbb{E}c(T,S)$, then $\mathsf{S}_{\mu}$ reduces to the Wasserstein correlation in (31).
5 Comparison of $\preccurlyeq_{ccx}$ with related dependence orderings
In this section, we compare the conditional convex order with well-established dependence orderings from the literature. While Theorem 1.6 reveals strong connections between the conditional convex order and the concordance order in terms of dimension-reduced bivariate SI random vectors, the two dependence orderings are fundamentally different; see Figure 3 and Table 1. As we show in Proposition 5.1, the two orderings coincide for SI random vectors. We then discuss related global dependence orderings studied in [22, 59, 60] and demonstrate that $\preccurlyeq_{ccx}$ provides a more natural and powerful framework for ordering the strength of functional dependence.
@@ 5.1 Comparison of $\preccurlyeq_{ccx}$ with the concordance order
where the lower orthant and upper orthant order on the right-hand side of (37) are defined by $P(Y\leq y,X\leq x)\leq P(Y^{\prime}\leq y,X^{\prime}\leq x)$ and $P(Y>y,X>x)\leq P(Y^{\prime}>y,X^{\prime}>x)$ for all $x,y\in\mathbb{R}$, respectively, see e.g. [32, Definition 2.4]. By definition, $\leq_{c}$ is a pure dependence order since $(Y^{\prime},X^{\prime})\leq_{c}(Y^{\prime},X^{\prime})$ implies $X\stackrel{{\scriptstyle\mathrm{d}}}{{=}}X^{\prime}$ and $Y\stackrel{{\scriptstyle\mathrm{d}}}{{=}}Y^{\prime}$. Note that for equal marginal distributions, the bivariate concordance order is equivalent to the lower (and also to the upper) orthant order; see Proposition B.2.
To relate the concordance order and the conditional convex order, let us consider for the moment bivariate random vectors $(V,U)$ and $(V^{\prime},U^{\prime})$ with $U(0,1)$ marginals. Then we are in the setting of bivariate copulas. Using the notation $\eta_{V|U}^{v}(t)=P(V\leq v\mid U=t)$ in (9) and disintegration, we observe that the concordance order can be characterized by
@@
An extension of the above reasoning to arbitrary marginals yields the following result. It states that $\preccurlyeq_{ccx}$ and $\leq_{c}$ coincide for bivariate SI random vectors from the same Fréchet class.
@@
Figure 3: Visualization of the concordance order (left panel) and the conditional convex order (right panel) for bivariate copulas. Here, $W$ and $M$ denote the lower and upper Fréchet copula and $\Pi$ the independence copula. While $W$ and $M$ are the uniquely determined minimal and maximal elements in concordance order, the conditional convex order admits the independence copula as a global minimal element and all perfectly dependent copulas as global maximal elements.
@@
Invariance under increasing transformations $(a(Y),X)\leq_{c}(a(Y^{\prime}),X^{\prime})$ $(a(Y),\mathbb{X})\preccurlyeq_{ccx}(a(Y^{\prime}),\mathbb{X}^{\prime})$
Invariance under decreasing transformations $(b(Y),X)\geq_{c}(b(Y^{\prime}),X^{\prime})$ $(b(Y),\mathbb{X})\preccurlyeq_{ccx}(b(Y^{\prime}),\mathbb{X}^{\prime})$
Invariance under bijective transformations $(Y,{\mathbf{h}}(\mathbb{X}))\preccurlyeq_{ccx}(Y^{\prime},\mathbf{h}(\mathbb{X}^{\prime}))$
Table 1: Basic properties of the bivariate concordance order and the conditional convex order, where $A$ is a closed subset of $[0,1]$, $a$ and $b$ are strictly increasing/decreasing functions, and $\mathbf{h}$ is a bijective function